The average breaking strength of steel rods is specified to be 18.5 thousand lbs. for this a sample of 14 rods was tested. The mean and standard deviation obtained were 17.85 and 1.955, respectively. Test the significance of the deviation by using 5% level of significance. (t=2.16 at 5%. d.f = 13) O


1
Expert's answer
2021-01-27T03:12:27-0500

Given, "n=14, \sigma=1.955, \bar{x}=17.85, \mu=18.5, \alpha=0.05"

Hypothesis is:

"H_O:\mu=18.5" Vs "H_1:\mu\ne18.5"



"\Z=(\bar{x}-\mu)/(\sigma| \sqrt n)=(17.85-18.5)/(1.955 | \sqrt 14)=-1.244"

P-value at 0.05 and "\Z=1.244" corresponds to 0.2135

Therefore, "2P(\Z" "\text{\textless}-1.244)=2\times 0.10749=0.2150"

"P-value=0.2150"

Since "P-value(0.2150) \gt 0.05" , we failed to reject "H_O"

In conclusion from the above, it is clear that the average breaking strength of steel rods is 18.5.